If an 8kg mass is pulled up an incline at a constant speed, determine
(a) the Tension in the string.
(b)the coefficient of friction between the 8kg mass and the surface (Assume there is no friction between the cord and the pulley).​
1
Expert's answer
2021-02-10T18:23:42-0500

a)

"T=mgcos \alpha=80cos \alpha~N,"

b)

"\mu=tan\alpha."